PTC is a substance that has a strong bitter taste for some people and is tasteless for others. The ability to taste PTC is inherited. About 75% of Italians can taste PTC, for example. You want to estimate the proportion of Americans with at least one Italian grandparent who can taste PTC.


Starting with the 75% estimate for Italians, how large a sample must you collect in order to estimate the proportion of PTC tasters within ±0.04 with 98 % confidence? (Give your answers as a whole number.)
Answer #1

Solution :

Given that,

= 0.75

1 - = 0.25

margin of error = E = 0.0

Z/2 = 2.326

sample size = n = (Z / 2 / E)2 * * (1 - )

= (2.326 / 0.04)2 * 0.75 * 0.25

= 634.02

sample size = n = 634